There are two plant nurseries in a small town. They are called Tumbleweed and Native Roots. If neither advertises, Tumbleweed makes $80,000 a month in profits and Native Roots makes $95,000. Advertising would cost each firm $20,000 a month. If only one firm advertises, that firm increases sales by $50,000 a month whereas the nonadvertising firm loses out. If Tumbleweed doesn't advertise but Native Roots does, Tumbleweed loses $30,000 a month. If Native Roots doesn't advertise but Tumbleweed does, it loses $35,000 a month. If both advertise, they increase revenue by $15,000 each. Insofar as they grow their products from the ground, they don't have any increased costs when they have increased sales (that is, their marginal cost of production is $0).

Part 1(2 pts)

What is the amount of profit Tumbleweed makes when both advertise?   $   

How much profit does Native Roots make when both advertise?   $   

Part 2(1 pt)

What outcome is predicted (that is, the Nash equilibrium) for these two firms, given the figures above?

Choose one:

A. Both firms advertise.
B. Tumbleweed advertises, but Native Roots doesn't.
C. Native Roots advertises, but Tumbleweed doesn't.
D. Neither firm advertises.
Answer #1

From the given information we can find out the profits of these two companies depending upon the 4 situations.

When both of the companies dont advertise their profits = 80k and 95k

when they both advertise there is an increase in sale for both companies by 15k but they have to incur the cost of advertisement that is 20k per month so if both of them advertise the profits will be 80+15-20=75k and 95+15-20=90k

now if Tumbleweed advertises but Native roots doesnot so there will be an 50k increase in the sales for Tumbleweed the profit of Tumbleweed will be 80+50-20 = 110k but there will be a loss of 35K for Natives roots thus the profit will be 95-35=60k

now if Native roots advertise the profit will be 95+50-20= 125k and Tumble weed doesnot advertise thus it looses 30k the profit of Tumbleweed will be 80-30=50k

thus the payoff matrix will look like

TW/NR A NA
A 75k,90k 110k,60k
NA 50k,125k 80k,95k

now for Tumbleweed Not advertising is strictly dominated by Advertising whatever be the strategy of native roots thus Tumble weed always prefer to advertise

so the new matrix will look like

TW/NR A NA
A 75k,90k 110k,60k

now for Native Roots not advertising strategy is strictly dominated by advertising strategy whatever be tumbleweed chooses . So Native Roots will always chhose to advertise thus here both the firm will advertise and NAsh equilibrium will be (A,A)